OpenStudy (anonymous):

find two consecutive negative odd integers such that the square of the lesser integer is 40 more than the square of the greater integer

OpenStudy (anonymous):

x^2 = (x + 2)^2 + 40 x^2 = x^2 + 4x + 4 + 40 x = -11 -11 and -9 121 - 81 = 40 correct -11 and -9